Your heart rate rises when you exercise. During aerobic exercise, your heart rate should be at least 0.7(220-a), but no more then 0.85(220-a), where a is your age. Find the target range if u are 14 years old.
Answer by checkley71(8403)

.7(220-14)=.7*206=144.2 for the lower limit.
.85(220-14)=.85*206=175.1 for the upper limit.